### v1

# Ansible Lightspeed: A Code Generation Service for IT Automation

Abstract.
The availability of Large Language Models (LLMs) which can generate code, has made it possible to create tools that improve developer productivity. Integrated development environments or IDEs which developers use to write software are often used as an interface to interact with LLMs. Although many such tools have been released, almost all of them focus on general-purpose programming languages. Domain-specific languages, such as those crucial for IT automation, have not received much attention. Ansible is one such YAML-based IT automation-specific language. Red Hat Ansible Lightspeed with IBM Watson Code Assistant, further referred to as Ansible Lightspeed, is an LLM-based service designed explicitly for natural language to Ansible code generation.
In this paper, we describe the design and implementation of the Ansible Lightspeed service and analyze feedback from thousands of real users. We examine diverse performance indicators, classified according to both immediate and extended utilization patterns along with user sentiments. The analysis shows that the user acceptance rate of Ansible Lightspeed suggestions is higher than comparable tools that are more general and not specific to a programming language. This remains true even after we use much more stringent criteria for what is considered an accepted model suggestion, discarding suggestions which were heavily edited after being accepted. The relatively high acceptance rate results in higher-than-expected user retention and generally positive user feedback. This paper provides insights on how a comparatively small, dedicated model performs on a domain-specific language and more importantly, how it is received by users.
Large Language Models, Generative Models, Code Completion, IDE, User Study, Ansible
copyright: none
1. Introduction
Ansible is a YAML (a plain-text data serialization language) based domain-specific language dedicated to IT automation. A typical Ansible project consists of YAML files, which are organized into playbooks (programs) and roles (libraries). Figure 1 shows an Ansible playbook with it’s sub-components. Playbooks consist of plays, which are a mapping between hosts and the tasks (sequential execution units) that run on those hosts.
Tasks contain a natural language description in the form of a name field, a module name defining the action to execute, and keys (or options) configuring the action.
Figure 1. A typical Ansible playbook structure.
Ansible Lightspeed (RedHat, 2023) is a generative AI service that utilizes the IBM Watson Code Assistant (WCA) to produce code recommendations based on Ansible best practices. WCA for Ansible is an extension of Ansible Wisdom (Pujar et al., 2023), which is a transformer decoder model of 350 million parameters trained on custom natural language, code and Ansible data.
An example of a user interaction with Ansible Lightspeed is presented Figure 2. With Ansible Lightspeed, you can build an Ansible playbook step by step, by providing a natural language description of an Ansible task using its name field. The description becomes a prompt for the model to generate the code of the task.
Code generation models, and some of the systems that utilize them, have emerged as powerful tools for software developers and system analysts (GitHub, 2021; Tab9, 2018; Replit, 2016; Amazon, 2023). Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Peng et al., 2023; Barke et al., 2023).
Pujar et al. (2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on Natural Language to Ansible task generation on benchmark data. But to our knowledge, there has been no published study that has involved thousands of users on the use of code assistants for domain-specific languages. Ansible Lightspeed is available since June  2023(RedHat, 2023), and thousands of developers have interacted with the system and consented to provide valuable feedback. In this paper we analyze the interactions and feedback of users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data.
We first analyze usage trends on a temporal basis, to understand when Ansible Lightspeed is used. Next we see if users continue to use Ansible Lightspeed and if so, how regularly. Most importantly, we check to what degree users accept suggestions made by Ansible Lightspeed. This acceptance rate is a crucial comparison metric with other code completion tools in the industry. It is also essential for us to understand how accurate the accepted suggestions were from the user’s perspective. Which is why we perform an edit analysis of the accepted suggestions and come up with a more stringent criteria for acceptance, where we consider a suggestion as accepted only if the user does not edit most or critical parts of the suggestion after accepting it. We analyze and share user feedback of the Ansible Lightspeed service. The key highlights of our analysis are the following:
  • We show that Lightspeed usage is the highest during working hours, on weekdays, suggesting that developers are probably using Ansible Lightspeed for work.
  • Using N-day user retention metric, we show that a Lightspeed user is 4 times more likely than an average Android app user and 3 times more likely than an iOS app user, to keep using Lightspeed after 30 days.
  • We are the first code completion tool to share user retention statistics. This can be used as a baseline for similar tools in the future.
  • We show that the acceptance rate of Ansible Lightspeed is higher than that of similar, but more general, IDE based code completion tools.
  • We develop a much more stringent acceptance criteria and show that even with this stringent criteria, Ansible Lightspeed’s acceptance rate is higher than the simple acceptance rate of other code completion tools.
  • A majority of the users give Ansible Lightspeed 4 and above rating on a scale of 1 to 5, with 5 being the highest rating.
With this analysis we demonstrate how domain expertise, knowledge of language utilization and the application of best practices for system design and user experience can significantly improve recommendation results and user acceptance rates. Given that we have a domain specific model, targeting a well defined use case, we are able to achieve these relatively high user acceptance rates with a relatively small model of size 350 million parameters.
The paper makes the following contributions:
  • A detailed description of the Ansible Lightspeed system and analysis framework (Section 2)
  • An analysis along multiple dimensions of the usage of a service by thousands of users (Section 3).
  • A much more accurate way to measure acceptance rate of model code suggestions by users, by accounting for user edits (Section 3.3).
(a) User writes the prompt.
(b) Ansible Lightspeed provides an inline suggestion.
(c) User accepts the suggestion by hitting the ‘Tab’ key.
Figure 2. Ansible Lightspeed’s workflow in the text editor.
2. Ansible Lightspeed
Figure 3. Ansible Lightspeed application architecture. A: VS Code extension, B: inference pipeline, C: content matching pipeline, D: analysis framework.
At a very high level, the Ansible Lightspeed platform consists of the following components:
  • A code editor extension that captures and transmits prompt and context information for inference and captures feedback from the user to improve model and service quality.
  • An inference pipeline for combining and processing natural language and Ansible-YAML to get code suggestions from a large language model.
  • A content matching pipeline that finds training examples that are similar to the code suggestions.
  • An analysis framework that collects and processes feedback data to make it consumable by a wide range of analysis tools, ultimately for the purpose of improving model quality and user experience.
Ansible Lightspeed is based on a client-server architecture, presented in Figure 3. The client is the VS Code Ansible extension which is published on the Visual Studio Marketplace (Community, 2021b) and provides a bulk of language specific features for Ansible. The server consists of an application service that handles authentication, authorization and entitlement checking, exposes APIs for inference, feedback and content matching, processing of inference prompts and model recommendations to improve quality. It is designed to be highly scalable and highly available, deployed in a multi-zone and multi-region configuration, ensuring that a large number of requests can be served simultaneously and with minimal latency.
2.1. Code Editor Extension
The Ansible extension relies on the Ansible Language server (Community, 2021a) that supports the Language server protocol (Microsoft, 2016) to provide language specific features like code auto-completion, diagnostics, hover and go-to features. The language service invokes Ansible Lint (Community, 2020) in the background to identify any issues with the Ansible file that the user is currently editing and provides real-time diagnostic information in the VS Code problems tab.
To use Ansible Lightspeed from within the code editor, the user must enable the Ansible Lightspeed feature in the Ansible extension settings, ensure the file type of the current file that’s being edited is ”Ansible” and then login to the Ansible Lightspeed service. After the login is successful, the user is ready to initiate requests and receive code suggestions from the Ansible Lightspeed service via the Ansible extension.
As shown in Figure 2, when a user starts writing Ansible code with a task description of the form - name: ¡Natural language prompt¿, and then presses enter, a completion request is sent to the Ansible Lightspeed service. The request includes all the content of the file up to the cursor position. After the server responds, the editor displays the response as greyed out text, which the user can either accept by pressing the tab key or reject by pressing the escape key.
For the accepted suggestions, the users can see the content matches in the training data by opening the Ansible panel within the editor. This sends a contentmatches request to the Ansible Lightspeed service which returns a list of content matches. This list of content matches allows the user to review the sources which possibly match the suggestion thus improving the user confidence in the accepted suggestion.
$2.$ Inference Pipeline
When the Ansible Lightspeed service receives a completion request from the client, it has to return a code suggestion. The completion request consists of a prompt field and the value is the Ansible code before the cursor position where the inline suggestion is triggered in the code editor. Within the service, the completion request processing is passed through three stages: pre-processing, inference and post-processing.
Pre-processing
In the pre-processing stage the prompt is checked for valid YAML syntax and the indentations are adjusted based on the requirements of the inference request. After that, it is passed through Ansible Anonymizer (Budder, 2023) to remove any personal, identifiable information. The cleaned up prompt is then passed to the next stage where the inference request is sent to the model server.
Inference
IBM Watson Code Assistant for Red Hat Ansible, or WCA-Ansible, is the model that powers Ansible Lightspeed.
WCA-Ansible is a transformer-based decoder trained from scratch on natural language, source code, and Ansible data. We use the Hugging Face Transformers framework (Wolf et al., 2020) to train a 350 million parameters model with a context window of size 1024; this model utilizes fewer parameters than the model available to subscription holders. The model is pre-trained with about 143 billion tokens and then fine-tuned with about 500k natural language to Ansible task samples. The fine-tuning data is created using a technique similar to the one of Ansible Wisdom (Pujar et al., 2023). The pre-training and fine-tuning data is license filtered to remove all data samples coming from files or repositories with restrictive licenses. We also filter the data to remove hate, abuse and profanity. The pre-training data is also deduplicated at a file level. The model server runs on machines containing A10101010 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
Post-processing
After the inference response is received by the Ansible Lightspeed service the response is again passed through the Ansible Anonymizer to cleanup any personal identifiable information that might be present in the model output. To ensure the suggestion follows the good practices of Ansible content development defined by the Ansible community (Community, 2019), the response is passed through Ansible Risk Insights or ARI tool (et al, 2020). The ARI tool runs multiple custom rules to modify the suggestion as required to ensure best practices such as invoking modules using their fully-qualified collection names and reusing Ansible variables defined in the context. It also redacts any credentials and removes and/or replaces any deprecated Ansible syntax from the suggestions. After this post-processing stage is complete, the response is sent back to the client.
2.3. Content Matching Pipeline
When the Ansible Lightspeed service receives a contentmatches request from the client with the accepted suggestion as a payload, it triggers the following actions.
The service encodes the recommendation using a Sentence-Transformers (SBERT (Reimers and Gurevych, 2019)) model. The service sends the encoded suggestion text to a search instance that utilizes the k-nearest neighbor (k-NN) algorithm (Duda et al., 1973) to return the top three nearest matches to the text suggestion. The search instance is indexed with the Ansible data used for fine-tuning the model.
Each match consists of information like the name and URL of the repository, the file path in the repository, the license, and the match score as per the k-NN algorithm.
2.4. Analysis Framework
Our data analysis framework uses an event tracking platform to monitor and record a wide range of events, including user interactions, system activities, and user feedback for Ansible Lightspeed. This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system. It includes user-initiated actions like completion requests, suggestion acceptance, and service feedback, as well as service-related events like server requests, matching sources, and post-processing logs.
The analysis framework consists of two parts: (1) a usage analyser evaluating engagement of the users, and (2) a suggestion analyzer evaluating the quality of the code suggestions.
Usage analyser
After collecting the telemetry data, we use event segmentation techniques to categorize and filter the data, allowing us to focus on specific event types and user behaviors. This helps us identify patterns, trends, and anomalies in the data, which in turn provides valuable insights into user preferences, satisfaction levels, and usage trends for Ansible Lightspeed.
Suggestion analyzer
Figure 4. Architecture of the suggestion analyser.
The second analysis is done on the user’s actions following the acceptance of the suggestions. We designed a Python framework to gather the playbook content of each user to study the modifications made to the accepted task suggestions. The architecture of the framework is shown in Figure 4. It runs in two steps: data processing and data analysis.
We begin processing data by sorting user data and only keeping essential information related to three specific event categories: Suggestion events (regarding suggestions provided by the model), User action events (indicating whether the user accepted or rejected the suggestion), and Content events (related to document content). After organizing the data into separate files for each user, we send it to the analyzer for a detailed analysis of each user’s data.
In the analyzer, we combine data from the sequential events explained above to create pairs of suggestion received and suggestion committed. Then, we break down the YAML content into lines and perform a sequence comparison. We use the SequenceMatcher class from the difflib library (Library, 2001) to compute the deltas. SequenceMatcher uses gestalt pattern matching (John, 1988), which aims to identify the longest contiguous matching sub-sequence without “junk” elements. This recursive process is applied to both sides of the matching sub-sequence within the sequences. Although it does not provide minimal edit sequences, it often aligns well with human perceptual expectations in sequence comparisons. Finally, we parse the YAML into objects to gain insights into the parts where the user has made edits.
This analysis provides us with two crucial insights: Firstly, it helps us determine if the user has entirely altered the suggestion, which is essential for assessing the actual acceptance of the suggestion. Secondly, it allows us to identify which parts of the suggestion have been modified, giving us an idea of how well the suggestion aligns with the user’s expectations.
3. Analysis
Our analysis aims to provide valuable insights into user behavior of the Ansible Lightspeed service and offer suggestions to improve user engagement and satisfaction. To achieve this, we address the following questions:
  • What are the usage trends of the service?
  • Do users continue to use the service after trying it?
  • Do users accept the suggestions proposed by the service and do they modify them?
  • How do users edit the code after accepting the suggestions?
  • How do users rate their experience using the explicit feedback mechanism?
By answering these questions, we can assess the service’s capacity to engage, satisfy, and retain users, as well as its potential for long-term influence. This multifaceted approach provides a comprehensive understanding of user behavior on the Ansible Lightspeed service and offers valuable insights for enhancing user experience and fostering continued engagement.
Our analysis is based on data collected from users who agreed to share their usage of the closed beta and free Tech Preview versions of the Ansible Lightspeed service from April 15, 2023 to February 4, 2024, a total of 296 days.
3.1. What are the usage trends of the service?
Table 1. Ansible Lightspeed user and usage trend.
Hourly Daily Weekly Monthly
Active user engagements 37 146 490 2.02k
Active users returning over time 35 123 349 796
New users 2 29 163 690
Total predictions 107 1.43k 6.07k 35.2k
Prediction requests from unique users 24 128 425 1.36k
Average prediction requests per user 4.46 11.2 14.3 25.9
Within the domain of evaluating service usability, the usage trend performance indicator plays a crucial role as a quantitative metric (Lewis and Sauro, 2021; Alzahrani et al., 2022). We analyze the total number of active users scrutinized at hourly, daily, weekly, and monthly intervals. Table 1 features varying engagement frequencies and diversification statistics, providing insights into user behavior and engagement patterns.
We are examining the total prediction requests viewed on hourly, daily, and weekly scales, and we notice several interesting patterns. One pattern shows that users tend to make more requests during work hours on weekdays. This increased activity during work hours highlights the service’s usefulness for work-related tasks. Another pattern reveals that some users have a more sporadic usage trend, particularly during weekends and evenings. This suggests that the service is also being used for non-work-related activities, such as personal projects or learning, indicating its versatility and potential to cater to a wide range of user needs and preferences.
The study of ‘Active users returning over time’ (second row of Table 1) provides us with an opportunity for a more in-depth examination of user retention. We expand upon this topic in the next subsection.
3.2. Do users continue to use the service after trying it?
Figure 5. N-Day User Retention Trend of Ansible Lightspeed. The dark green represents the percentage of users returning on Day N.
User retention metrics are crucial in app and service analytics, providing valuable insights into user engagement and app longevity (Aspinall et al., 2001). Our study utilizes the N-day User Retention metric to evaluate the Ansible Lightspeed service (Jarvis, 2023). This metric calculates the percentage of users who continue interacting with an app N days after their initial installation or first use, offering a vital assessment of the app’s ability to retain users over time. A high N-day retention percentage indicates that users find lasting value in the app’s features, content, or services, serving as a barometer of user satisfaction, the effectiveness of on-boarding processes, and the overall quality of the app. Figure 5 shows the N-day user retention of Ansible Lightspeed service.
To gain a better understanding of the N-day user retention, ideally we would have liked to compare Ansible Lightspeed with other code completion tools or IDE Plugins. However since N-day user retention is not available for similar tools or for VS Code plugins, we compare it with global data compiled by Business of Apps (Tafradzhiyski, 2023) for Android and iOS apps, as well as their user retention data for 2022 (Handa, 2022) which represents the worldwide average retention rate across 31 mobile app categories. Table 2 presents the results. On Day 1, the average retention rate is 25.3%, gradually decreasing to 5.7% by Day 30. When looking at platform-specific data, iOS has a Day 1 retention rate of 25.65%, while its Day 30 retention rate drops to 4.13%. Android, on the other hand, exhibits a Day 1 retention rate of 23.01%, with its Day 30 retention standing at 2.59%.
These statistics highlight the typical retention patterns in the iOS and Android ecosystems. Retention rates strongly decline over time. However, the Ansible Lightspeed service stands out as an exception to these trends. With a Day 1 retention rate of 44.79%, Lightspeed significantly outperforms global averages and even surpasses the iOS and Android platforms. Additionally, on Day 30, Lightspeed maintains a strong retention rate of 13.66%, demonstrating its ability to retain users over an extended period.
Table 2. Comparison of N-Day user retention of Ansible Lightspeed service with Android and iOS app trends.
Apps 1-Day 3-Day 7-Day 30-Day
Android apps 23.44% 11.30% 7.16% 2.96%
iOS apps 27.06% 13.76% 9.46% 4.40%
Ansible Lightspeed 44.79% 34.17% 33.68% 13.66%
Our findings reveal that Ansible Lightspeed would rank among the top Android applications. Specifically, Ansible Lightspeed falls within the range of the top 100 Android applications regarding user retention data (Chen, 2018). We are the first code completion tool to release N-Day user retention figures and we hope these figures act as baseline for other tools in the future.
$3.$ Do users accept the suggestions proposed by the service and do they modify them?
(a) Distribution of suggestion modifications in the accepted suggestions.
(b) Distribution of user edit patterns in the scenarios where they have made minor edits in the accepted suggestions.
Figure 6. Ansible Lightspeed suggestion modification and edits analysis.
The Ansible Lightspeed service provides inline suggestions in the editor generated from the task name and document content. Users receive these suggestions, which are always multi-line, after entering the task name and moving to the next line. Then, users can either accept the suggestion by pressing the ‘Tab’ key or reject it by pressing the ‘Esc’ key. This interaction is recorded as a telemetry event, and the data is used to evaluate the initial acceptance rate of the model’s suggestions.
We selected a cohort of 3,474 users from a pool of about 6k new users to conduct our user acceptance rate analysis. In the context of our study, we use the term, ”returning user” for any user who used the Ansible Lightspeed service for two or more days. We found 3,474 such returning users, and they have collectively made 74,583 requests to the Ansible Lightspeed service for task suggestions.
Our analysis reveals that out of 74,583 suggestions Lightspeed made, 36,259 suggestions were accepted by users, resulting in an initial acceptance rate of 48.62%.
Table 3. Summary of the user edit analysis.
Events Numbers
Total users analyzed 3,474
Total suggestion requests 74,583
Initially accepted suggestions 36,259
Fully accepted suggestions (0% edits) 21,642
Minor edits (¡ 50%) made after acceptance 5,247
Major edits ($\geq$ 50%) made after acceptance 2,641
Deleted suggestions after acceptance 6,433
Module changed in the accepted suggestion 296
To further evaluate these accepted requests, we examined the changes made by users to the accepted suggestions. Figure 6 and Table 3 analyze the edits made to the accepted suggestions.
In 21,642 instances (29.02% of the suggestions and 59.69% of the accepted ones), Ansible Lightspeed’s suggestions perfectly aligned with user requirements. The service accurately predicted the appropriate Ansible module, incorporating the correct options and sub-options, seemingly matching the users’ objectives. Users did not make any modifications to the endorsed suggestion, indicating a high level of satisfaction with the service’s recommendations.
In 5,247 scenarios (7.04% of the suggestions and 14.47% of the accepted ones), users made small adjustments (less than 50% edits) to the accepted suggestions. While Ansible Lightspeed accurately predicted the correct usage of the desired module, users modified one or more module keys (options) and/or values to better align with their specific needs. We examine the edits in greater detail later in the section.
In 2,641 of the remaining suggestions (3.54% of the suggestions and 7.28% of the accepted ones), we observe substantial modifications (50% edits or more) after acceptance, highlighting the crucial need for significant changes in the generated code to meet users’ requirements.
We noticed that users deleted 6,433 accepted suggestions after acceptance, which is 17.74% of the total initially accepted suggestions. Two possible explanations for this behaviour could be:
  1. (1)
    Users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to accept the suggestion for a thorough examination only to realize it was inaccurate, which made them delete the suggestion.
  2. (2)
    Users experimenting with the prompt or examining the deterministic nature of the model, resulting in multiple successive requests after discarding the received suggestion, to see if the answers vary.
The module holds significant importance in an Ansible task. Our analysis reveals that when less than 50% of edits were made, there were 296 instances where users edited the module after accepting the suggestion. This pattern underscores situations where the model’s suggestion did not align with users’ specific module requirements.
So if we consider as a form of rejection the cases where the module is changed or more than 50% of the suggestion is edited, or where the suggestions were deleted after they were first accepted, the actual acceptance rate reduces to 36.05% (down from the initial 48.62%).
We compared our acceptance rate with the acceptance rates of Google Machine Learning enabled code completion system (Maxim Tabachnyk, 2022) (MLECC) which is an internal tool at Google, Github CoPilot (Ziegler et al., 2022) and CodeCompose (Murali et al., 2023), which is an internal tool at Meta. These were the only other code completion tools we found to have used acceptance rate as a metric. Tab. 4 presents the comparison. Google MLECC (Maxim Tabachnyk, 2022) provides acceptance rates for single-line and multi-line code completion, for about 10k internal users, for 8 programming languages over 3 months. They provide user counts for both single-line and multi-line code completion, but don’t breakdown acceptance rate based on languages.
Ziegler et al. (2022) show Co-pilot acceptance rate for survey participants and for all users, for multiple programming languages. Acceptance rate for survey participants are always higher compared to all users, which is why we only show survey participant acceptance rates in Tab. 4. A total of 17,420 participated in their survey. They don’t share total users for each language.
Murali et al. (2023) show CodeCompose acceptance rates and user counts overall and split across multiple languages. We picked top 3 languages with highest acceptance rates. Ansible Lightspeed generates multi-line Ansible code suggestions.
Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and after discarding suggestions which had more than 50% edits.
Table 4. Comparison of acceptance rate of Ansible Lightspeed with the other code completion tools. Google MLECC (Maxim Tabachnyk, 2022) stands for ML enabled code completion. Ansible Lightspeed Initial accept are percent suggestions which were accepted by users. Edit¡50% are percent suggestions which were accepted and had less than 50% user edits, based on difflib SequenceMatcher(Library, 2001).
Service Acceptance Rates User Count
Google MLECC
single-line $C$ 25% 10k+
multi-line $C$ 34% 5k+
Github Co-Pilot 17,420
Typescript 26.3%
Javascript 31.4%
Python 30.8%
Other 23.4%
CodeCompose
Hack 22.5% 5.5k
Python 22.0% 10.7k
C 21.3% 201
All 22% 16k
Ansible Lightspeed
Initial accept (Ansible) 48.6% 3,474
Edit¡50% (Ansible) 36.1% 3,474
Finally, to better understand the users behavior, we also study which part of the code the users are editing when they do only small modification. Our findings are as follows:
  1. (1)
    1,503 times out of the 5,247 suggestions with minor edits, users made changes in both the “KEY” and the ”VALUE” parts of the suggested options of the module, added a pair option and value, or removed one. This kind of edit either changes the behavior of the task while staying within the context of the suggested module.
  2. (2)
    In 3,730 instances of minor edits, users opted to change only the “VALUE” component of the received suggestions. This involves personalized modifications to variable names, numeric values, and strings, aligning the suggestions precisely with their unique requirements.
  3. (3)
    We found only 14 such cases in the minor edits category, where the users edited only the ”KEY” component of the received task suggestion. Typically, these occurrences occur when a module has an option with multiple aliases, and the user decides to update an option with its alternative name.
3.4. How do users rate their experience using the explicit feedback mechanism?
Figure 7. Ansible Lightspeed Feedback User Interface.
In VS Code Editor, we have created a dedicated user interface (UI) in the sidebar for Ansible Lightspeed feedback, as shown in Figure 7. Within this UI, users can provide ratings on a scale of 1 to 5, along with written feedback on their experience using the service. We have received a total of 605 user ratings. The user sentiment analysis is performed on this data.
Star rating
Figure 8. Ansible Lightspeed’s user sentiments.
The rating in term of stars is presented Figure 8. The data reveals a high level of user satisfaction with the service and recommendations from the Ansible Lightspeed service. Notably, 56.8% of users express considerable satisfaction with the task recommendations, as indicated by their assignment of 4 and 5-star ratings.
However, the sentiment landscape becomes more nuanced when we examine lower rating categories. While the 3-star average rating of 15.9% indicates a moderate level of satisfaction, it’s important to note that users within this category find some value in the recommendations, representing a positive aspect. They recognize areas for improvement but are only somewhat dissatisfied. Our analysis also reveals a 27.3% average rating distribution encompassing 2-star and 1-star ratings. These ratings signify dissatisfaction with the suggestions. Although negative, their feedback serves as constructive input for refining and enhancing the model.
User feedback
In addition to the star rating, we did a thorough analysis of the textual users’ feedback for the Ansible Lightspeed service. We made a manual topic extraction and classification of the of the negative feedback (1 or 2 stars) and positive feedback (3 or 4 stars). The results are presented in Figures 9 and 10.
To provide an impression of the comments, we present a selection of feedback obtained from Ansible Lightspeed users. These quotes have been rephrased for clarity and anonymity, removing any references to internal details.
Negative feedback
Figure 9. Distribution of negative feedback.
The majority of the negative feedback, 66.49%, is coming from users that do not manage to get Ansible Lightspeed to work. These problems stem from various sources, including installation difficulties, authentication troubles, network problems, and bugs.
The second largest source of complaints, accounting for 15.71%, is due to poor suggestions provided by the model. This is often caused by the use of uncommon modules, as illustrated by the following example of feedback:
“I mainly use a specific collection that may not be part of training so suggestion were rarely helpful”
Users are also disappointed with the suggestions provided by the model because they are requesting tasks that fall outside of its current scope. For instance, they may ask for multi-task or playbook generation, but the model is only capable of generating single tasks:
“Has really bad recommendations. Tried creating a playbook with name ”Get …”” and it get me a module recommendation”
Finally, 5.79% of users are dissatisfied with their overall experience, specifically citing the need to re-authenticate periodically as an issue. The remaining 12.04% of comments are not informative.
We’re working to address user concerns by continuously improving the service’s documentation and robustness. We’re also expanding the model’s capabilities to handle more use cases and providing customization options so users can choose their preferred collections or provide their own.
Positive feedback
Figure 10. Distribution of positive feedback.
On the positive side, 19.72% of customers praise the service’s user-friendliness, highlighting its simplicity in setup and use. For instance, some users have mentioned the login and on-boarding process as being easy and fast. Additionally, a sizable 33.68% of favorable feedback emphasizes the service’s exceptional suggestion accuracy, which practically translates to recommendations for exact code production that align with users’ automation goals. Users have commented that the model not only recommends the correct module but also effectively utilizes module options and sub-options. Furthermore, 42.70% of users consider improved productivity as Ansible Lightspeed’s key benefit. This increase in productivity is supported by observable results like time savings, accelerated documentation activities, and improved process effectiveness.
To illustrate the positive sentiments from the users, we present a selection their feedback.
After the release of the Ansible Lightspeed service, multiple demo sessions were conducted for users. This is what one of the users has to say after using the service for the first time:
“The session was fantastic and was incredibly straightforward to follow. Although I’ve been coding for over two decades, Ansible is entirely new territory for me. Despite my lack of familiarity with Ansible, the AI made it remarkably user-friendly. Thanks to the service, I could create a playbook without any prior knowledge of Ansible, and I mean absolutely none.”
In this case, the user was new to Ansible, but they had a delightful experience writing their first playbook with the help of Ansible Lightspeed. Several remarks that also came from seasoned playbook developers indicate the usefulness of using Ansible Lightspeed in their daily workflow because of its efficiency.
“Lightspeed is genuinely amazing and has significantly impacted my productivity. It saves me from the hassle of constantly switching between Ansible documentation websites for reference.”
“This extension is quite handy for speeding up my daily work in the office. It provides helpful task suggestions, and the inline module suggestions really cut down on time spent.”
We also found out that upon using the service, users experimented with the prompts and found it to return the correct suggestions most of the time. This is what one of the users has to say about this:
“This was beyond my expectation that the tool could manage complex syntax without any issues. Also, the NLP capabilities are impressive. I tested it with different writing styles for the same prompts, and it consistently provided the expected results. For example, it handled both of these cases perfectly:
$11$ - name: Can we see the value of win_credential_results? $22$ - name: Print win_credential_results”
In addition to these remarks, we also observed interesting side effects, such as increased best practices in playbook writing when developers use Ansible Lightspeed. As Ansible Lightspeed runs Ansible Risk Insights (et al, 2020) and Ansible Lint as post-processing tasks, it ensures that the task suggestions follow the best practices. This is evident in the anecdote mentioned below:
“Ansible Lightspeed is fantastic because I don’t need to constantly remember recommended practices, such as using fully qualified collection names (fqcn) every time. The suggestions it provides align with these best practices. The suggestions help me write cleaner loops and maintain consistent white spaces, making my work much smoother.”
This underscores the significance of the user experience for AI-assisted code generation with Ansible Lightspeed. While the service is seen as valuable for playbook developers in IT automation, some feedback compared it with the sophistication of GPT-4. We are actively working to enhance Ansible Lightspeed for an improved experience for all Ansible playbook developers.
4. Related Work
LLMs have shown remarkable ability to generate code, with many recent models like GPT-4(Achiam et al., 2023), Llama(Touvron et al., 2023), StarCoder(Li et al., 2023) and others performing very well on code evaluation benchmarks like HumanEval(Chen et al., 2021) and MBPP(Austin et al., 2021).
A lot of work has also been done on developing coding assistants for IDEs based on LLMs. These would include tools that are internal to a company and are only accessible by internal users and those that are openly available. Murali et al. (2023) analyze CodeCompose, an internal code completion tool at Meta which is based on InCoder LLM (Fried et al., 2022), and show that the acceptance rate is about 22% across 9 programming languages for approximately 16k users. Similarly Maxim Tabachnyk (2022) analyze a code completion tool at Google and measure an acceptance rate of 25-34% over 10k+ Google-internal developers. Svyatkovskiy et al. (2020) introduce and evaluate IntelliCode Compose on multiple programming languages, but they use evaluation based on edit-distance compared to the ground truth and do not mention any user acceptance metrics or the total number of users. None of these works mention user sentiment or user retention over an extended period of time.
The openly available coding assistants would be GitHub Copilot (GitHub, 2021), Tab9 (Tab9, 2018), Replit (Replit, 2016), Amazon CodeWhisperer (Amazon, 2023) and Ansible Lightspeed (RedHat, 2023), among others. Among these, GitHub Copilot has been widely studied since it’s release. Nguyen and Nadi (2022) test Copilot on 33 LeetCode questions in four programming languages. Vaithilingam et al. (2022) perform a more user centred evaluation with 24 users to see how programmers use and perceive Copilot. Ziegler et al. (2022) perform an in-depth study of user acceptance, similar to what we do, but for multiple programming languages and show that the acceptance rate of Copilot suggestions for different user categories, for different programming languages, is approximately 20%-30%. Peng et al. (2023) study the impact of Co-pilot on the speed of programmers and find that AI pair programmers are 55.8% faster in implementing an HTTP server in JavaScript. However they do not provide information about acceptance rate of AI pair programmers. Yetiştiren et al. (2023) perform a comparative study of GitHub Copilot, Amazon CodeWhisperer and ChatGPT (OpenAI, 2022) (sibling model of InstructGPT å(Ouyang et al., 2022)) in terms of code quality metrics like code correctness, code security, code reliability and code maintainability but do not consider acceptance of code by real users. A lot of studies on coding assistants do not measure their performance in terms of acceptance rate. Among those that do, we found that Ansible Lightspeed has the highest initial acceptance rate of 48.6%.
We did not find any user retention figures for any of the existing code completion systems or for the VS Code plugins in the VS Code marketplace. In the absence of more relevant baselines, we compare Lightspeed retention rate with that of Android and iOS apps.
5. Limitations
While this analysis provides comprehensive insights into the impact of Ansible Lightspeed on user experience, there are certain considerations and limitations to be acknowledged.
We do not perform acceptance rate and edit analysis on all Ansible Lightspeed users. Our Telemetry data includes events from users that may not be Ansible programmers and simply want to try out the system; event data from these users would bias the acceptance rate and edit analysis. To mitigate this issue, we created a cohort of returning users, who used Ansible Lightspeed on two or more days.
Ansible Lightspeed’s acceptance rate is compared with models which are much more general, can be used for multiple programming languages and are likely bigger in terms of number of parameters and the amount of data used for training. These general models also tend to have a higher number of users overall and their performance on Ansible-YAML has not been studied or reported. One of the main purposes of our study was to explore the performance of a domain, language specific model vis-a-vis a more general model, with real users. The total number of users, 3,474, is sizable enough for us to be confident about our analysis.
Additionally, our evaluation of Ansible Lightspeed’s user retention lacked official data for comparisons with other code generation services. To address this gap, we compared its user retention with Android and iOS mobile apps, providing insights into its sustained user engagement. While the comparison methodology may not perfectly align with the domain, it highlights Ansible Lightspeed’s ability to retain users over time.
Despite these considerations, the analysis and insights presented here offer a substantial understanding of Ansible Lightspeed’s effectiveness, user engagement, and potential areas for improvement. Based on this work, we hope to continue investigations into domain-specific code generation tools and their impact on IT automation.
6. Conclusion and Future work
Pujar et al. (2023) show that a relatively small model fine-tuned on a specific Ansible task generation use case can outperform a much larger model on a benchmark dataset. In this paper we show that the suggestions made by such a specifically fine-tuned model can have relatively higher acceptance rate among a large set of users. We also show that the relatively higher acceptance rate results in better user retention compared to the average for Android and iOS apps and generally positive user feedback.
Currently we support only Ansible task generation, however we are working to expand the capabilities of the model to be able to write entire Ansible playbook based on user prompt. We also plan to develop capabilities for Ansible code explanation, debugging and as well as customization for specific users.
Acknowledgements.
We want to thank the following people for their help, support, and consultation while envisioning, building and deploying Ansible Lightspeed at Red Hat: Mattew Jones, Xavier Lecauchois, Brian King, Robin Bobbit, Scott Harwell, Craig Brandt, James Wong, Marty Turner, Goneri Le Bouder, Tami Takamiya, Sumit Jaiswal. We would also like to thank the model development team and the Ansible Risks Insights team at IBM Research as well as the Watson Code Assistant team.

### v2

@@
-Ansible Lightspeed: A Code Generation Service for IT Automation
+Insights from the Usage of the Ansible Lightspeed Code Completion Service
 Abstract.
@@
 Although many such tools have been released, almost all of them focus on general-purpose programming languages.
-Domain-specific languages, such as those crucial for IT automation, have not received much attention.
+Domain-specific languages, such as those crucial for Information Technology (IT) automation, have not received much attention.
 Ansible is one such YAML-based IT automation-specific language.
-Red Hat Ansible Lightspeed with IBM Watson Code Assistant, further referred to as Ansible Lightspeed, is an LLM-based service designed explicitly for natural language to Ansible code generation.
-In this paper, we describe the design and implementation of the Ansible Lightspeed service and analyze feedback from thousands of real users.
-We examine diverse performance indicators, classified according to both immediate and extended utilization patterns along with user sentiments.
-The analysis shows that the user acceptance rate of Ansible Lightspeed suggestions is higher than comparable tools that are more general and not specific to a programming language.
-This remains true even after we use much more stringent criteria for what is considered an accepted model suggestion, discarding suggestions which were heavily edited after being accepted.
-The relatively high acceptance rate results in higher-than-expected user retention and generally positive user feedback.
-This paper provides insights on how a comparatively small, dedicated model performs on a domain-specific language and more importantly, how it is received by users.
+Ansible Lightspeed is an LLM-based service designed explicitly to generate Ansible YAML given natural language prompt.
+This paper first presents the design and implementation of the Ansible Lightspeed service.
+We then evaluate its utility to developers using diverse indicators, including extended utilization, analysis of user rejected suggestions, as well as analysis of user sentiments.
+The analysis is based on data collected for 10,696 real users including 3,910 returning users.
+The code for Ansible Lightspeed service and the analysis framework is made available for others to use.
+To our knowledge, our study is the first to involve thousands of users in evaluating code assistants for domain-specific languages.
+We propose an improved version of user acceptance rate and we are the first code completion tool to present N-Day user retention figures.
+With our findings we provide insights into the effectiveness of small, dedicated models in a domain-specific context.
+We hope this work serves as a reference for software engineering and machine learning researchers exploring code completion services for domain-specific languages in particular and programming languages in general.
 Large Language Models, Generative Models, Code Completion, IDE, User Study, Ansible
-copyright: none
+$c$s: Computing methodologies Natural language processingccs: Software and its engineering Context specific languagesccs: Software and its engineering Application specific development environments
 1. Introduction
-Ansible is a YAML (a plain-text data serialization language) based domain-specific language dedicated to IT automation.
-A typical Ansible project consists of YAML files, which are organized into playbooks (programs) and roles (libraries).
+Figure 1.
+A typical Ansible playbook structure consists of playbook definition and tasks list.
+Ansible Lightspeed generates a task, given the task name and the preceding context.
+Module name is the first term generated by Lightspeed, followed by the module associated keys and values.
+[…]
+Ansible (Ansible Community, 2024) is a YAML based domain-specific language dedicated to IT automation.
+It is one of the most widely used infrastructure as code tools (Guerriero et al., 2019; StackExchange, 2023).
+Since it is open-source, we do not know the precise number of developers who use Ansible but we can get an indication of its popularity with the Ansible python community package (Team, 2024b) getting over 5 million downloads per month (Flynn, 2024).
+Ansible also has an active community on GitHub (Team, 2024a) with about 5,000 contributors and about 54,000 commits on a yearly basis, which is likely a small fraction of the actual user base.
+[^1] Another indication of Ansible’s popularity is from Ansible’s social media accounts of Reddit (Team, 2024c) and X (Team, 2024d), counting 63,000 readers and 72,000 followers, respectively.
+A typical Ansible project is organized into playbooks (programs) and roles (libraries).
 Figure 1 shows an Ansible playbook with it’s sub-components.
@@
 Tasks contain a natural language description in the form of a name field, a module name defining the action to execute, and keys (or options) configuring the action.
-Figure 1. A typical Ansible playbook structure.
-Ansible Lightspeed (RedHat, 2023) is a generative AI service that utilizes the IBM Watson Code Assistant (WCA) to produce code recommendations based on Ansible best practices.
-WCA for Ansible is an extension of Ansible Wisdom (Pujar et al., 2023), which is a transformer decoder model of 350 million parameters trained on custom natural language, code and Ansible data.
-An example of a user interaction with Ansible Lightspeed is presented Figure 2.
-With Ansible Lightspeed, you can build an Ansible playbook step by step, by providing a natural language description of an Ansible task using its name field.
-The description becomes a prompt for the model to generate the code of the task.
 Code generation models, and some of the systems that utilize them, have emerged as powerful tools for software developers and system analysts (GitHub, 2021; Tab9, 2018; Replit, 2016; Amazon, 2023).
-Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Peng et al., 2023; Barke et al., 2023).
+Studies show that AI tools for general-purpose programming languages improve productivity and also suggest that novice programmers may benefit more from such tools (Weber et al., 2024; Peng et al., 2023; Barke et al., 2023).
+However, generic models do not perform as well on domain-specific languages.
 Pujar et al.
-(2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on Natural Language to Ansible task generation on benchmark data.
-But to our knowledge, there has been no published study that has involved thousands of users on the use of code assistants for domain-specific languages.
-Ansible Lightspeed is available since June  2023(RedHat, 2023), and thousands of developers have interacted with the system and consented to provide valuable feedback.
-In this paper we analyze the interactions and feedback of users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data.
-We first analyze usage trends on a temporal basis, to understand when Ansible Lightspeed is used.
-Next we see if users continue to use Ansible Lightspeed and if so, how regularly.
+(2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on benchmark Natural Language to Ansible-task generation test data.
+But this work does not involve real users in their development environment interacting with a model.
+Red Hat Ansible Lightspeed (RedHat, 2023), further referred to as Ansible Lightspeed or simply Lightspeed, is a generative AI service that utilizes watsonx Code Assistant for Ansible (WCA), an extension of Ansible Wisdom (Pujar et al., 2023), to produce code recommendations based on best practices.
+An example of a user interaction with Ansible Lightspeed is presented in Figure 2.
+With Ansible Lightspeed, you can build an Ansible playbook step-by-step, by providing a natural language description of an Ansible task using its name field.
+The description becomes a prompt for the model to generate the code of the task, which usually consists of the module name followed by module keys and values.
+(a) User writes the prompt.
+(b) Ansible Lightspeed provides an inline suggestion.
+(c) User accepts the suggestion by hitting the ‘Tab’ key.
+[…]
+Figure 2.
+Ansible Lightspeed’s workflow in the text editor.
+Users receive Lightspeed suggestions, which are almost always multi-line, after entering the task name and moving to the next line.
+Then, users can either accept the suggestion by pressing the ‘Tab’ key or reject it by pressing the ‘Esc’ key.
+Ansible Lightspeed was available since April 15 2023 to a limited number of users as part of a closed beta release, and then generally to everyone as part of a free Tech Preview version from June 2023  (RedHat, 2023) until May 15 2024.
+During this period thousands of developers interacted with the system.
+In this paper we analyze the interactions and feedback of 10,696 users of the closed beta and free Tech Preview versions of Ansible Lightspeed, who consented to share their data.
+To our knowledge, there has been no published study, with several thousand users, on the use of code assistants for domain-specific languages.
+As a part of our study we analyze usage trends on a temporal basis, to understand when Ansible Lightspeed is used.
+We see if users continue to use Ansible Lightspeed and if so, how regularly.
 Most importantly, we check to what degree users accept suggestions made by Ansible Lightspeed.
@@
 It is also essential for us to understand how accurate the accepted suggestions were from the user’s perspective.
-Which is why we perform an edit analysis of the accepted suggestions and come up with a more stringent criteria for acceptance, where we consider a suggestion as accepted only if the user does not edit most or critical parts of the suggestion after accepting it.
-We analyze and share user feedback of the Ansible Lightspeed service.
-The key highlights of our analysis are the following:
-  • We show that Lightspeed usage is the highest during working hours, on weekdays, suggesting that developers are probably using Ansible Lightspeed for work.
-  • Using N-day user retention metric, we show that a Lightspeed user is 4 times more likely than an average Android app user and 3 times more likely than an iOS app user, to keep using Lightspeed after 30 days.
+This is done through an edit analysis of the accepted suggestions.
+The edit analysis tells us how the user edited the model suggestion before using it.
+This gives valuable insights on whether the model suggestion was truly useful and clues on how we can improve it.
+As a result of edit analysis, we also get a more effective and stringent criteria for acceptance, where we consider a suggestion as accepted only if the user does not edit most or critical parts of the suggestion after accepting it.
+Finally, we analyze and share user qualitative feedback of the Ansible Lightspeed service.
+The paper makes the following contributions:
+  • A detailed description of the Ansible Lightspeed system and analysis framework.
+The code for Ansible lightspeed [^2] https://github.com/ansible/ansible-ai-connect-service and the analysis framework [^3] https://github.com/ansible-community/lightspeed-analysis-framework has been made available on GitHub.
   • We are the first code completion tool to share user retention statistics. This can be used as a baseline for similar tools in the future.
-  • We show that the acceptance rate of Ansible Lightspeed is higher than that of similar, but more general, IDE based code completion tools.
-  • We develop a much more stringent acceptance criteria and show that even with this stringent criteria, Ansible Lightspeed’s acceptance rate is higher than the simple acceptance rate of other code completion tools.
-  • A majority of the users give Ansible Lightspeed 4 and above rating on a scale of 1 to 5, with 5 being the highest rating.
-With this analysis we demonstrate how domain expertise, knowledge of language utilization and the application of best practices for system design and user experience can significantly improve recommendation results and user acceptance rates.
-Given that we have a domain specific model, targeting a well defined use case, we are able to achieve these relatively high user acceptance rates with a relatively small model of size 350 million parameters.
-The paper makes the following contributions:
-  • A detailed description of the Ansible Lightspeed system and analysis framework (Section 2)
-  • An analysis along multiple dimensions of the usage of a service by thousands of users (Section 3).
-  • A much more accurate way to measure acceptance rate of model code suggestions by users, by accounting for user edits (Section 3.3).
-(a) User writes the prompt.
-(b) Ansible Lightspeed provides an inline suggestion.
-(c) User accepts the suggestion by hitting the ‘Tab’ key.
-Figure 2. Ansible Lightspeed’s workflow in the text editor.
+  • We propose an improved acceptance criteria that checks if the accepted suggestion was actually used by the user.
+  • We show that Ansible Lightspeed’s acceptance rate is higher than that of other, more general, code completion tools.
+Given that we have a domain-specific model, targeting a well defined use case, we are able to achieve high user acceptance rates with a relatively small model of size 350 million parameters.
+Our analysis of user made edits to model suggestions, presents insights on how users use the suggestion and how we can improve it.
+This study also provides a reference point for software engineering and machine learning researchers investigating the adoption of code completion services, especially for domain-specific languages.
 2. Ansible Lightspeed
-Figure 3. Ansible Lightspeed application architecture. A: VS Code extension, B: inference pipeline, C: content matching pipeline, D: analysis framework.
+[…]
 At a very high level, the Ansible Lightspeed platform consists of the following components:
@@
   • An inference pipeline for combining and processing natural language and Ansible-YAML to get code suggestions from a large language model.
-  • A content matching pipeline that finds training examples that are similar to the code suggestions.
+  • A content matching pipeline for attribution, that finds training examples that are similar to the code suggestions and provides the location of the corresponding repositories.
   • An analysis framework that collects and processes feedback data to make it consumable by a wide range of analysis tools, ultimately for the purpose of improving model quality and user experience.
@@
 This list of content matches allows the user to review the sources which possibly match the suggestion thus improving the user confidence in the accepted suggestion.
-$2.$ Inference Pipeline
+2.2. Inference Pipeline
 When the Ansible Lightspeed service receives a completion request from the client, it has to return a code suggestion.
@@
 In the pre-processing stage the prompt is checked for valid YAML syntax and the indentations are adjusted based on the requirements of the inference request.
-After that, it is passed through Ansible Anonymizer (Budder, 2023) to remove any personal, identifiable information.
+After that, it is passed through Ansible Anonymizer (Bouder, 2023) to remove any personal, identifiable information.
 The cleaned up prompt is then passed to the next stage where the inference request is sent to the model server.
 Inference
-IBM Watson Code Assistant for Red Hat Ansible, or WCA-Ansible, is the model that powers Ansible Lightspeed.
+IBM watsonx Code Assistant for Red Hat Ansible, or WCA-Ansible, is the model that powers Ansible Lightspeed.
 WCA-Ansible is a transformer-based decoder trained from scratch on natural language, source code, and Ansible data.
@@
 The pre-training data is also deduplicated at a file level.
-The model server runs on machines containing A10101010 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
+The model server runs on machines containing A100100100100 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
 Post-processing
@@
 Each match consists of information like the name and URL of the repository, the file path in the repository, the license, and the match score as per the k-NN algorithm.
-2.4. Analysis Framework
-Our data analysis framework uses an event tracking platform to monitor and record a wide range of events, including user interactions, system activities, and user feedback for Ansible Lightspeed.
+2.4. Data Collection and Analysis Framework
+Figure 4.
+Ansible Lightspeed Feedback User Interface.
+The users can rate their experience using emojis, which we call star-rating.
+The right most emoji would indicate the best experience or a 5-star rating, and the left most emoji would indicate the worst experience or a 1-star rating.
+Users are also expected to write a few words to explain their rating in ”Tell us why?” section.
+[…]
+Ansible Lightspeed collects explicit user feedback, and implicit event tracking.
+For the explicit feedback, we have created a dedicated user interface (UI) in the sidebar of VS Code Editor as shown in Figure 4.
+Within this UI, users can provide ratings on a scale of 1 to 5 stars, represented as emojis with rightmost emoji being 5 stars (the best) and leftmost emoji being 1 star (the worst).
+Along with the rating the users also provide a written feedback on their experience using the service.
+More on this explicit feedback in Sec.
+4.
+An event tracking platform monitors and records a wide range of events about user interactions, and system activities.
 This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system.
-It includes user-initiated actions like completion requests, suggestion acceptance, and service feedback, as well as service-related events like server requests, matching sources, and post-processing logs.
-The analysis framework consists of two parts: (1) a usage analyser evaluating engagement of the users, and (2) a suggestion analyzer evaluating the quality of the code suggestions.
+It includes user-initiated actions like completion requests, suggestion acceptance, as well as service-related events like server requests, matching sources, and post-processing logs.
+The analysis framework consists of two parts: (1) a usage analyzer evaluating engagement of the users, and (2) a suggestion analyzer evaluating the quality of the code suggestions.
 Usage analyser
@@
 Figure 5. Architecture of the suggestion analyser.
+[…]
 The second analysis is done on the user’s actions following the acceptance of the suggestions.
@@
 Secondly, it allows us to identify which parts of the suggestion have been modified, giving us an idea of how well the suggestion aligns with the user’s expectations.
-3. Analysis
-Our analysis aims to provide valuable insights into user behavior of the Ansible Lightspeed service and offer suggestions to improve user engagement and satisfaction.
-To achieve this, we address the following questions:
-  • What are the usage trends of the service?
-  • Do users continue to use the service after trying it?
-  • Do users accept the suggestions proposed by the service and do they modify them?
-  • How do users edit the code after accepting the suggestions?
-  • How do users rate their experience using the explicit feedback mechanism?
-By answering these questions, we can assess the service’s capacity to engage, satisfy, and retain users, as well as its potential for long-term influence.
-This multifaceted approach provides a comprehensive understanding of user behavior on the Ansible Lightspeed service and offers valuable insights for enhancing user experience and fostering continued engagement.
-Our analysis is based on data collected from users who agreed to share their usage of the closed beta and free Tech Preview versions of the Ansible Lightspeed service from April 15, 2023 to February 4, 2024, a total of 296 days.
-3.1. What are the usage trends of the service?
-Table 1. Ansible Lightspeed user and usage trend.
-Hourly Daily Weekly Monthly
-Active user engagements 37 146 490 2.02k
-Active users returning over time 35 123 349 796
-New users 2 29 163 690
-Total predictions 107 1.43k 6.07k 35.2k
-Prediction requests from unique users 24 128 425 1.36k
-Average prediction requests per user 4.46 11.2 14.3 25.9
-Within the domain of evaluating service usability, the usage trend performance indicator plays a crucial role as a quantitative metric (Lewis and Sauro, 2021; Alzahrani et al., 2022).
-We analyze the total number of active users scrutinized at hourly, daily, weekly, and monthly intervals.
-Table 1 features varying engagement frequencies and diversification statistics, providing insights into user behavior and engagement patterns.
-We are examining the total prediction requests viewed on hourly, daily, and weekly scales, and we notice several interesting patterns.
-One pattern shows that users tend to make more requests during work hours on weekdays.
-This increased activity during work hours highlights the service’s usefulness for work-related tasks.
-Another pattern reveals that some users have a more sporadic usage trend, particularly during weekends and evenings.
-This suggests that the service is also being used for non-work-related activities, such as personal projects or learning, indicating its versatility and potential to cater to a wide range of user needs and preferences.
-The study of ‘Active users returning over time’ (second row of Table 1) provides us with an opportunity for a more in-depth examination of user retention.
-We expand upon this topic in the next subsection.
-3.2. Do users continue to use the service after trying it?
+3. Analysis of User Interactions
+The goal of our analysis is to determine if the users find the service useful and how we can improve the service to better fit user needs. To achieve this, we try to address the following questions:
+  • When do users use Ansible Lightspeed?
+  • Do users continue to use the service after trying it once?
+  • Do users modify accepted suggestions?
+  • At what rate do users accept suggestions?
+  • Why do users edit module names in accepted suggestions?
+Our analysis is based on data collected from users who agreed to share their usage data of the closed beta and free Tech Preview version of the Ansible Lightspeed service from April 15, 2023, to May 15, 2024, a total of 397 days.
+Figure 6.
+Number of Ansible Lightspeed completion requests per day.
+Weekdays are displayed in green [ ] and weekends in red [ ].
+Average completion requests per day - Monday: 832.69, Tuesday: 1065.08, Wednesday: 990.92, Thursday: 1063.19, Friday: 864.77, Saturday: 241.25, Sunday: 264.90.
+The data is normalized based on different time zones from where the users made the completion requests.
+[…]
+3.1. Temporal Analysis
+Lightspeed’s goal is to improve the productivity of developers who use Ansible for work.
+Understanding when requests were made could give some hint that Lightspeed is being used for work related queries.
+Figure 6 shows all Lightspeed completion requests [^4] The figures here represent telemetry events, which may contain duplicates.
+The duplicates will inflate the request count but we believe the trends highlighted below will be remain the same.
+on a daily basis for the entire analysis period.
+Requests on weekdays are shown in green and those on weekends are shown in red.
+One observation is that users tend to make more requests on an average weekday than on an average weekend.
+We found that ‘Tuesdays’ had the maximum average completion requests per day (average of 1065.08 completion requests) and ‘Saturdays’ had the lowest average completion requests per day (average of 241.25 completion requests).
+The reason for this weekday and weekend difference could be that a significant chunk of our users are using Lightspeed for work, during weekdays, rather than for study or side projects.
+However, we need more analysis to be sure.
+In Figure 6 we can also see a significant decline in the completion requests between December 23, 2023, and January 12, 2024.
+This could be due to the Holiday season when many offices are usually closed.
+Q: When do users use Ansible Lightspeed?
+Users make more requests during weekdays, with noticeable drops on weekends and around holidays, which could mean that many users are using Lightspeed for work related requests.
+3.2. User Retention
 Figure 7. N-Day User Retention Trend of Ansible Lightspeed. The dark green represents the percentage of users returning on Day N.
-User retention metrics are crucial in app and service analytics, providing valuable insights into user engagement and app longevity (Aspinall et al., 2001).
+[…]
+User retention metrics are crucial in-app and service analytics, providing valuable insights into user engagement and app longevity (Aspinall et al., 2001).
 Our study utilizes the N-day User Retention metric to evaluate the Ansible Lightspeed service (Jarvis, 2023).
-This metric calculates the percentage of users who continue interacting with an app N days after their initial installation or first use, offering a vital assessment of the app’s ability to retain users over time.
-A high N-day retention percentage indicates that users find lasting value in the app’s features, content, or services, serving as a barometer of user satisfaction, the effectiveness of on-boarding processes, and the overall quality of the app.
+This metric calculates the percentage of users who continue interacting with an app N days after their initial installation or first use.
+Because of this metric’s utility in serving as a barometer for user satisfaction, it is used in many industries, including by iOS and Android app developers (Tafradzhiyski, 2023).
+For instance, iOS apps have an average Day 30 retention of 4.13%, while the same number for Android apps is 2.59% (Handa, 2022).
 Figure 7 shows the N-day user retention of Ansible Lightspeed service.
-To gain a better understanding of the N-day user retention, ideally we would have liked to compare Ansible Lightspeed with other code completion tools or IDE Plugins.
-However since N-day user retention is not available for similar tools or for VS Code plugins, we compare it with global data compiled by Business of Apps (Tafradzhiyski, 2023) for Android and iOS apps, as well as their user retention data for 2022 (Handa, 2022) which represents the worldwide average retention rate across 31 mobile app categories.
-Table 2 presents the results.
-On Day 1, the average retention rate is 25.3%, gradually decreasing to 5.7% by Day 30.
-When looking at platform-specific data, iOS has a Day 1 retention rate of 25.65%, while its Day 30 retention rate drops to 4.13%.
-Android, on the other hand, exhibits a Day 1 retention rate of 23.01%, with its Day 30 retention standing at 2.59%.
-These statistics highlight the typical retention patterns in the iOS and Android ecosystems.
-Retention rates strongly decline over time.
-However, the Ansible Lightspeed service stands out as an exception to these trends.
-With a Day 1 retention rate of 44.79%, Lightspeed significantly outperforms global averages and even surpasses the iOS and Android platforms.
-Additionally, on Day 30, Lightspeed maintains a strong retention rate of 13.66%, demonstrating its ability to retain users over an extended period.
-Table 2. Comparison of N-Day user retention of Ansible Lightspeed service with Android and iOS app trends.
-Apps 1-Day 3-Day 7-Day 30-Day
-Android apps 23.44% 11.30% 7.16% 2.96%
-iOS apps 27.06% 13.76% 9.46% 4.40%
-Ansible Lightspeed 44.79% 34.17% 33.68% 13.66%
-Our findings reveal that Ansible Lightspeed would rank among the top Android applications.
-Specifically, Ansible Lightspeed falls within the range of the top 100 Android applications regarding user retention data (Chen, 2018).
-We are the first code completion tool to release N-Day user retention figures and we hope these figures act as baseline for other tools in the future.
-$3.$ Do users accept the suggestions proposed by the service and do they modify them?
-(a) Distribution of suggestion modifications in the accepted suggestions.
+As can be seen, of all the users who use the service on their Day 0, 44.79% return on Day 1, 34.17% on Day 3 and so on.
+13.66% users who first used the service 30 days prior, continue to use the service on Day 30.
+Based on N-day user retention, we define a returning user, as someone who returned to use the service at least once more after Day 0.
+Out of a total of about 10,696 unique users, 3,910 or 36.6% were returning users.
+Since returning users are much more engaged with the service, we assume that they may be using the service for real work, rather than testing it using toy examples.
+All further analysis is on this cohort of 3,910 are returning users.
+Q: Do users continue to use the service after trying it once?
+On Day 1 the average retention rate is 44.79%.
+On Day 30 it is 13.66%.
+3,910 users out of total 10,696 users, or about 36.5%, are returning users.
+3.3. Edit Analysis
+(a) Distribution of suggestion modifications in the accepted suggestions. ”Module Changed” indicates Minor edit suggestions in which module was changed.
 (b) Distribution of user edit patterns in the scenarios where they have made minor edits in the accepted suggestions.
 Figure 8. Ansible Lightspeed suggestion modification and edits analysis.
-The Ansible Lightspeed service provides inline suggestions in the editor generated from the task name and document content.
-Users receive these suggestions, which are always multi-line, after entering the task name and moving to the next line.
-Then, users can either accept the suggestion by pressing the ‘Tab’ key or reject it by pressing the ‘Esc’ key.
-This interaction is recorded as a telemetry event, and the data is used to evaluate the initial acceptance rate of the model’s suggestions.
-We selected a cohort of 3,474 users from a pool of about 6k new users to conduct our user acceptance rate analysis.
-In the context of our study, we use the term, ”returning user” for any user who used the Ansible Lightspeed service for two or more days.
-We found 3,474 such returning users, and they have collectively made 74,583 requests to the Ansible Lightspeed service for task suggestions.
-Our analysis reveals that out of 74,583 suggestions Lightspeed made, 36,259 suggestions were accepted by users, resulting in an initial acceptance rate of 48.62%.
-Table 3. Summary of the user edit analysis.
+[…]
+The interactions shown in Figure 2 are recorded as telemetry events, [^5] There are duplicate telemetry events.
+We can’t be sure why there are duplicate events, but it may be as simple as network issues or user hitting the enter key twice, instead of once.
+We filter out duplicates and consider only unique requests for our analysis.
+and this telemetry data is used to calculate the initial acceptance rate of Lightspeed’s suggestions.
+We use the data of 3,910 returning users for our user acceptance rate analysis, since we consider returning users to be more engaged.
+These returning users made a total of 62,099 completion requests to Ansible Lightspeed and received that many suggestions.
+Out of 62,099 suggestions received by returning users, 40,938 suggestions were accepted, resulting in an initial acceptance rate of 65.92%.
+Table 1. Summary of the user edit analysis. A returning user is a user who engaged with the system for more than 1 day. All numbers below, except Total users, are for returning users.
 Events Numbers
-Total users analyzed 3,474
-Total suggestion requests 74,583
-Initially accepted suggestions 36,259
-Fully accepted suggestions (0% edits) 21,642
-Minor edits (¡ 50%) made after acceptance 5,247
-Major edits ($\geq$ 50%) made after acceptance 2,641
-Deleted suggestions after acceptance 6,433
-Module changed in the accepted suggestion 296
-To further evaluate these accepted requests, we examined the changes made by users to the accepted suggestions. Figure 6 and Table 3 analyze the edits made to the accepted suggestions.
-In 21,642 instances (29.02% of the suggestions and 59.69% of the accepted ones), Ansible Lightspeed’s suggestions perfectly aligned with user requirements.
-The service accurately predicted the appropriate Ansible module, incorporating the correct options and sub-options, seemingly matching the users’ objectives.
-Users did not make any modifications to the endorsed suggestion, indicating a high level of satisfaction with the service’s recommendations.
-In 5,247 scenarios (7.04% of the suggestions and 14.47% of the accepted ones), users made small adjustments (less than 50% edits) to the accepted suggestions.
-While Ansible Lightspeed accurately predicted the correct usage of the desired module, users modified one or more module keys (options) and/or values to better align with their specific needs.
-We examine the edits in greater detail later in the section.
-In 2,641 of the remaining suggestions (3.54% of the suggestions and 7.28% of the accepted ones), we observe substantial modifications (50% edits or more) after acceptance, highlighting the crucial need for significant changes in the generated code to meet users’ requirements.
-We noticed that users deleted 6,433 accepted suggestions after acceptance, which is 17.74% of the total initially accepted suggestions. Two possible explanations for this behaviour could be:
+Total users 10,696
+Returning users 3,910
+Total suggestions 62,099
+Initially accepted suggestions 40,938
+Fully accepted suggestions (0% edits) 24,811
+Minor edits (¡ 50%) made after acceptance 5,672
+Major edits ($\geq$ 50%) made after acceptance 2,713
+Deleted suggestions after acceptance 7,436
+Module changed, but not major edit 306
+To understand if these accepted suggestions were used by users, we further examined if changes were made by users to the accepted suggestions.
+Table 1 and Figure 8.a gives a high level overview of the edits made by returning users on accepted suggestions.
+In 24,811 instances (60.60% of the accepted suggestions), Ansible Lightspeed’s suggestions were used without edits so we can assume that they aligned with user requirements.
+In 5,672 scenarios (13.85% of the accepted suggestions), users made small adjustments, less than 50% edits [^6] Suggestion analyzer part of Sec.
+2.4, which explains how edits are calculated., to the accepted suggestions.
+An example of such a light or minor edit would be, when Ansible Lightspeed accurately predicts the correct module, users modified one or more module keys (options) and/or values to better align with their specific needs.
+In 2,713 instances (6.62% of the accepted suggestions), we observe substantial modifications (50% edits or more) after acceptance, indicating that while users accepted the suggestion, they had to heavily modify it before they could actually use it.
+We also noticed that users deleted 7,436 suggestions after acceptance, which is 18.16% of the total initially accepted suggestions.
+Currently there is no way to tell why the users would delete accepted suggestions and move on to next request.
+One explanation is that users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to initially accept and then delete after closer review.
+Another possible reason could be that they were experimenting with different prompts trying one after another and discarding the answers.
+We further analyzed 5,672 instances which were lightly edited, to make sure that they were indeed minor edits. Figure 8.b gives a break down of different kinds of minor edits.
+The first observation was that there were 306 instances of minor edits, or 5.4% of all accepted suggestions, in which the module name was also edited.
+A module could not be considered a minor edit because the keys and values which immediately follow depend on the type of module.
+If the model gets the module wrong, it is more likely to get rest of the suggestion wrong as well.
+If the module name is edited, we consider the suggestion to be rejected.
+We found that in a majority of remaining instances, the users opted to change only the ‘VALUE’ component of the received suggestions.
+This involves personalized modifications to variable names, numeric values, and strings, something which maybe hard for Lightspeed to predict based only on the name field prompt and preceding context.
+In a much smaller proportion users made changes in both the ‘KEY’ and the ‘VALUE’ parts of the suggested options of the module, added a pair option and value, or removed one.
+This kind of edit slightly changes the behavior of the task while staying within the context of the suggested module.
+Instances in which only the ’KEY’ was changed were negligible.
+Q: Do users modify accepted suggestions? Yes users modify accepted suggestions. 18.16% of initially accepted suggestions were deleted. 6.62% initially accepted suggestions were heavily modified.
+3.4. User Acceptance Rate
+The edit analysis described above shows that the acceptance rate as currently defined can be misleading.
+Which is why we define a new acceptance rate metric, called Edit¡50% acceptance rate.
+Edit¡50% Acceptance rate is the percentage of accepted suggestions, in which we consider a suggestion accepted only if the following conditions are fulfilled:
   1. (1)
-    Users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to accept the suggestion for a thorough examination only to realize it was inaccurate, which made them delete the suggestion.
+    The suggestion is initially accepted by the user.
   2. (2)
-    Users experimenting with the prompt or examining the deterministic nature of the model, resulting in multiple successive requests after discarding the received suggestion, to see if the answers vary.
-The module holds significant importance in an Ansible task.
-Our analysis reveals that when less than 50% of edits were made, there were 296 instances where users edited the module after accepting the suggestion.
-This pattern underscores situations where the model’s suggestion did not align with users’ specific module requirements.
-So if we consider as a form of rejection the cases where the module is changed or more than 50% of the suggestion is edited, or where the suggestions were deleted after they were first accepted, the actual acceptance rate reduces to 36.05% (down from the initial 48.62%).
-We compared our acceptance rate with the acceptance rates of Google Machine Learning enabled code completion system (Maxim Tabachnyk, 2022) (MLECC) which is an internal tool at Google, Github CoPilot (Ziegler et al., 2022) and CodeCompose (Murali et al., 2023), which is an internal tool at Meta.
+    After acceptance, the suggestion is not deleted.
+  3. (3)
+    Suggestion has minimal edits or less than 50% of the suggestion is edited.
+  4. (4)
+    Critical components, like the module name in case of Ansible, should not be edited.
+As per this definition, the Edit¡50% acceptance rate for Ansible Lightspeed comes out to 49.08%, compared to the initial acceptance rate of 65.92%.
+We feel Edit¡50% acceptance rate more closely aligns with users perception of what they consider accepted, compared to the initial acceptance rate.
+Table.
+2 shows Ansible Lightspeed acceptance rate compared to other IDE based code completion tools.
+We compared our acceptance rate with the initial acceptance rates of Google Machine Learning enabled code completion system (Maxim Tabachnyk, 2022) (MLECC) which is an internal tool at Google, Github CoPilot (Ziegler et al., 2022) and CodeCompose (Murali et al., 2023), which is an internal tool at Meta.
 These were the only other code completion tools we found to have used acceptance rate as a metric.
-Tab.
-4 presents the comparison.
-Google MLECC (Maxim Tabachnyk, 2022) provides acceptance rates for single-line and multi-line code completion, for about 10k internal users, for 8 programming languages over 3 months.
-They provide user counts for both single-line and multi-line code completion, but don’t breakdown acceptance rate based on languages.
+Google MLECC (Maxim Tabachnyk, 2022) provides acceptance rates for single-line and multi-line code completion, but don’t breakdown acceptance rate based on languages.
 Ziegler et al.
-(2022) show Co-pilot acceptance rate for survey participants and for all users, for multiple programming languages.
-Acceptance rate for survey participants are always higher compared to all users, which is why we only show survey participant acceptance rates in Tab.
-4.
-A total of 17,420 participated in their survey.
-They don’t share total users for each language.
-Murali et al.
-(2023) show CodeCompose acceptance rates and user counts overall and split across multiple languages.
-We picked top 3 languages with highest acceptance rates.
-Ansible Lightspeed generates multi-line Ansible code suggestions.
-Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and after discarding suggestions which had more than 50% edits.
+(2022) for GitHub Co-pilot and Murali et al.
+(2023) for CodeCompose share acceptance rate for individual languages.
+We show the top three languages with best acceptance rate.
+The acceptance rate for these three baselines is what what we call the initial acceptance rate.
+It merely tells us whether the user accepted the suggestion offered without telling us how the user actually used the suggestion.
+It does not give any insight into whether the user accepted the suggestion as is, or edited the suggestion or altogether deleted it.
+Ansible Lightspeed generates multi-line Ansible code suggestions. Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and Edit¡50% acceptance rate.
 Table 2.
@@
 Google MLECC
-single-line $C$ 25% 10k+
-multi-line $C$ 34% 5k+
+single-line CC 25% 10k+
+multi-line CC 34% 5k+
 Github Co-Pilot 17,420
@@
 Ansible Lightspeed
-Initial accept (Ansible) 48.6% 3,474
-Edit¡50% (Ansible) 36.1% 3,474
-Finally, to better understand the users behavior, we also study which part of the code the users are editing when they do only small modification. Our findings are as follows:
-  1. (1)
-    1,503 times out of the 5,247 suggestions with minor edits, users made changes in both the “KEY” and the ”VALUE” parts of the suggested options of the module, added a pair option and value, or removed one.
-This kind of edit either changes the behavior of the task while staying within the context of the suggested module.
-  2. (2)
-    In 3,730 instances of minor edits, users opted to change only the “VALUE” component of the received suggestions.
-This involves personalized modifications to variable names, numeric values, and strings, aligning the suggestions precisely with their unique requirements.
-  3. (3)
-    We found only 14 such cases in the minor edits category, where the users edited only the ”KEY” component of the received task suggestion.
-Typically, these occurrences occur when a module has an option with multiple aliases, and the user decides to update an option with its alternative name.
-3.4. How do users rate their experience using the explicit feedback mechanism?
-Figure 7. Ansible Lightspeed Feedback User Interface.
-In VS Code Editor, we have created a dedicated user interface (UI) in the sidebar for Ansible Lightspeed feedback, as shown in Figure 7.
-Within this UI, users can provide ratings on a scale of 1 to 5, along with written feedback on their experience using the service.
-We have received a total of 605 user ratings.
-The user sentiment analysis is performed on this data.
-Star rating
+Initial accept (Ansible) 65.9% 3,910
+Edit¡50% (Ansible) 49.1% 3,910
+Q: At what rate do users accept suggestions?
+Initial acceptance rate is misleading since users can delete or heavily modify suggestions.
+65.9% of suggestions are initially accepted.
+We define new acceptance rate, Edit¡50% Acceptance rate.
+49.1% of model suggestions were accepted as per Edit¡50% Acceptance rate.
+Ansible Lightspeed has relatively higher acceptance for Ansible, compared to more general tools for various programming languages.
+3.5. Rejected Suggestion Analysis
+There are a total of 1,710 initially accepted suggestions which were explicitly edited by the users at the module level.
+We consider such suggestions to be rejected for Edit¡50% Acceptance rate.
+However, the users must have found something useful in the suggestion for them to accept it and then spend sometime editing it.
+We analyse these edits in order to better understand how the model suggestions differ from user expectations.
+3.5.1. User preference
+In order to remove ambiguity in module names, Ansible Lint encourages users to use Fully Qualified Collection Name, or FQCN (Sbarnea, 2023), when declaring modules.
+In keeping with this best practice, Ansible Lightspeed always generates module names in FQCN format.
+On the other hand it is quite possible that some users prefer the shorter version of the module name.
+For example, the best practice FQCN reference would be ansible.builtin.debug while some users may prefer to refer to the same module as debug.
+We find that in 346, or 20.2% of, instances the users edited the model suggestion and changed the FQCN module name to it’s shortened format.
+The body of the generated Ansible task often had little to no edits, indicating that the FQCN module name, which is correct, is not what they prefer.
+3.5.2. YAML reorganization
+As per our observation, in the process of writing an Ansible playbook users often reorganize the YAML structure.
+This reorganization involves adding new valid Ansible keys like block, tag, register, loop and become.
+The model generated suggestion is generally part of the reorganized playbook, either in the edited or original format, and is likely to have inspired the reorganization.
+Such edits indicate that the generated suggestion is not necessarily incorrect, but is definitely incomplete and can be better organized.
+There are at least 211 (12.3%) such instances of reorganization.
+3.5.3. Functionally similar modules
+Module confusion usually occurs when there are multiple functionally similar modules which can be used to implement user’s intent and the user is at least somewhat familiar with one of them.
+One example of this is command and shell modules, which are often interchangeable.
+In addition, functionalities of many other modules can be implemented by command or shell modules.
+It is possible that users are not aware of other modules, or they may prefer another module than the one that lightspeed generated.
+Generated shell commands are also likely to be incorrect, which could inspire more user edits.
+There are 352 edits (20.6%) which are related to the command/shell module.
+Apart from command or shell module related edits, we found 336 (19.6%) instances of edits related to modules which have similar functionalities and can be used interchangeably.
+There are another 586 (28%) instances where the user edited the generated module. Understanding these edits requires further analysis.
+Q: Why do users edit module names in accepted suggestions?
+Suggestion is edited to remove FQCN based on user preference.
+YAML is reorganized to add more details which the model suggestion is missing.
+Certain functionalities can be implemented in multiple ways.
+Users change the suggestion based on their preferred implementation.
+4. User Feedback
+A total of 605 users have provided rating and comments using the the sidebar for Ansible Lightspeed in the VS Code Editor, as shown in Figure 4.
+4.1. Star rating
 Figure 9. Ansible Lightspeed’s user sentiments.
-The rating in term of stars is presented Figure 8.
-The data reveals a high level of user satisfaction with the service and recommendations from the Ansible Lightspeed service.
-Notably, 56.8% of users express considerable satisfaction with the task recommendations, as indicated by their assignment of 4 and 5-star ratings.
-However, the sentiment landscape becomes more nuanced when we examine lower rating categories.
-While the 3-star average rating of 15.9% indicates a moderate level of satisfaction, it’s important to note that users within this category find some value in the recommendations, representing a positive aspect.
-They recognize areas for improvement but are only somewhat dissatisfied.
-Our analysis also reveals a 27.3% average rating distribution encompassing 2-star and 1-star ratings.
-These ratings signify dissatisfaction with the suggestions.
-Although negative, their feedback serves as constructive input for refining and enhancing the model.
-User feedback
-In addition to the star rating, we did a thorough analysis of the textual users’ feedback for the Ansible Lightspeed service.
-We made a manual topic extraction and classification of the of the negative feedback (1 or 2 stars) and positive feedback (3 or 4 stars).
-The results are presented in Figures 9 and 10.
-To provide an impression of the comments, we present a selection of feedback obtained from Ansible Lightspeed users.
-These quotes have been rephrased for clarity and anonymity, removing any references to internal details.
+[…]
+A breakdown of ratings in terms of stars is presented in Figure 9.
+Notably, a total of 57% of users express considerable satisfaction with the service, as indicated by their assignment of 4 and 5-star ratings.
+The 3-star rating by 15.8% of users indicates a moderate level of satisfaction.
+In this case users do find some value in using Ansible Lightspeed, but expect more improvement.
+About 27.2% users appeared dissatisfied and gave a rating of 2 and 1-star.
+4.2. User Comments
+Apart from star-rating, users also submit feedback in the form of textual comments.
+By analyzing these comments we hope to find areas for improvement and also know what users like about the service.
+We manually review and classify negative feedback (1 or 2 stars) and positive feedback (4 or 5 stars) into various reasons.
+We also share some examples of real user comments, both positive and negative.
+These quotes have been rephrased for clarity and anonymity, removing any references to user details.
 Negative feedback
 Figure 10. Distribution of negative feedback.
+[…]
+The breakdown of negative comments based on manual review is presented in Figure 10.
 The majority of the negative feedback, 66.49%, is coming from users that do not manage to get Ansible Lightspeed to work.
@@
 The second largest source of complaints, accounting for 15.71%, is due to poor suggestions provided by the model.
-This is often caused by the use of uncommon modules, as illustrated by the following example of feedback:
+This could sometimes be caused by the use of uncommon modules, as illustrated by the following example of feedback:
 “I mainly use a specific collection that may not be part of training so suggestion were rarely helpful”
-Users are also disappointed with the suggestions provided by the model because they are requesting tasks that fall outside of its current scope.
-For instance, they may ask for multi-task or playbook generation, but the model is only capable of generating single tasks:
+Some users are disappointed with the suggestions provided by the model because their requests fall outside of Lightspeed’s current scope.
+For instance, they may ask for multi-task or playbook generation, but the model is currently only trained for generating single tasks:
 “Has really bad recommendations. Tried creating a playbook with name ”Get …”” and it get me a module recommendation”
 Finally, 5.79% of users are dissatisfied with their overall experience, specifically citing the need to re-authenticate periodically as an issue. The remaining 12.04% of comments are not informative.
-We’re working to address user concerns by continuously improving the service’s documentation and robustness.
-We’re also expanding the model’s capabilities to handle more use cases and providing customization options so users can choose their preferred collections or provide their own.
 Positive feedback
 Figure 11. Distribution of positive feedback.
-On the positive side, 19.72% of customers praise the service’s user-friendliness, highlighting its simplicity in setup and use.
-For instance, some users have mentioned the login and on-boarding process as being easy and fast.
-Additionally, a sizable 33.68% of favorable feedback emphasizes the service’s exceptional suggestion accuracy, which practically translates to recommendations for exact code production that align with users’ automation goals.
-Users have commented that the model not only recommends the correct module but also effectively utilizes module options and sub-options.
-Furthermore, 42.70% of users consider improved productivity as Ansible Lightspeed’s key benefit.
-This increase in productivity is supported by observable results like time savings, accelerated documentation activities, and improved process effectiveness.
-To illustrate the positive sentiments from the users, we present a selection their feedback.
-After the release of the Ansible Lightspeed service, multiple demo sessions were conducted for users. This is what one of the users has to say after using the service for the first time:
-“The session was fantastic and was incredibly straightforward to follow.
-Although I’ve been coding for over two decades, Ansible is entirely new territory for me.
-Despite my lack of familiarity with Ansible, the AI made it remarkably user-friendly.
-Thanks to the service, I could create a playbook without any prior knowledge of Ansible, and I mean absolutely none.”
-In this case, the user was new to Ansible, but they had a delightful experience writing their first playbook with the help of Ansible Lightspeed.
-Several remarks that also came from seasoned playbook developers indicate the usefulness of using Ansible Lightspeed in their daily workflow because of its efficiency.
+[…]
+Breakdown of reasons for positive experience is presented in Figure 11.
+Most users, about 42.7%, consider improved productivity as Ansible Lightspeed’s key benefit.
+Multiple users mentioned that they noticeably saved time by using Lightspeed.
 “Lightspeed is genuinely amazing and has significantly impacted my productivity. It saves me from the hassle of constantly switching between Ansible documentation websites for reference.”
 “This extension is quite handy for speeding up my daily work in the office. It provides helpful task suggestions, and the inline module suggestions really cut down on time spent.”
-We also found out that upon using the service, users experimented with the prompts and found it to return the correct suggestions most of the time. This is what one of the users has to say about this:
+33.7% of users found suggestion accuracy to be the most impressive part of Ansible Lightspeed.
+Users have commented that the model not only recommends the correct module but also the associated module options and values.
 “This was beyond my expectation that the tool could manage complex syntax without any issues.
@@
 $11$ - name: Can we see the value of win_credential_results? $22$ - name: Print win_credential_results”
-In addition to these remarks, we also observed interesting side effects, such as increased best practices in playbook writing when developers use Ansible Lightspeed.
-As Ansible Lightspeed runs Ansible Risk Insights (et al, 2020) and Ansible Lint as post-processing tasks, it ensures that the task suggestions follow the best practices.
-This is evident in the anecdote mentioned below:
-“Ansible Lightspeed is fantastic because I don’t need to constantly remember recommended practices, such as using fully qualified collection names (fqcn) every time.
-The suggestions it provides align with these best practices.
-The suggestions help me write cleaner loops and maintain consistent white spaces, making my work much smoother.”
-This underscores the significance of the user experience for AI-assisted code generation with Ansible Lightspeed.
-While the service is seen as valuable for playbook developers in IT automation, some feedback compared it with the sophistication of GPT-4.
-We are actively working to enhance Ansible Lightspeed for an improved experience for all Ansible playbook developers.
+About 19.7% of customers praise the service’s user-friendliness, highlighting its simplicity in setup and use.
+For instance, some users have mentioned the login and on-boarding process as being easy and fast.
+Remaining 3.9% of the users said best part was the general experience of using the tool.
 5. Related Work
@@
 (2023) analyze CodeCompose, an internal code completion tool at Meta which is based on InCoder LLM (Fried et al., 2022), and show that the acceptance rate is about 22% across 9 programming languages for approximately 16k users.
-Similarly Maxim Tabachnyk (2022) analyze a code completion tool at Google and measure an acceptance rate of 25-34% over 10k+ Google-internal developers.
+Similarly, Maxim Tabachnyk (2022) analyze a code completion tool at Google and measure an acceptance rate of 25-34% over 10k+ Google-internal developers.
 Svyatkovskiy et al.
@@
 None of these works mention user sentiment or user retention over an extended period of time.
-The openly available coding assistants would be GitHub Copilot (GitHub, 2021), Tab9 (Tab9, 2018), Replit (Replit, 2016), Amazon CodeWhisperer (Amazon, 2023) and Ansible Lightspeed (RedHat, 2023), among others.
-Among these, GitHub Copilot has been widely studied since it’s release.
+The openly available coding assistants would be GitHub Copilot (GitHub, 2021), Tab9 (Tab9, 2018), Replit (Replit, 2016) and Amazon CodeWhisperer (Amazon, 2023) among others.
+Among these, GitHub Copilot has been widely studied since its release.
 Nguyen and Nadi (2022) test Copilot on 33 LeetCode questions in four programming languages.
 Vaithilingam et al.
-(2022) perform a more user centred evaluation with 24 users to see how programmers use and perceive Copilot.
+(2022) perform a more user-centred evaluation with 24 users to see how programmers use and perceive Copilot.
 Ziegler et al.
@@
 (2023) study the impact of Co-pilot on the speed of programmers and find that AI pair programmers are 55.8% faster in implementing an HTTP server in JavaScript.
-However they do not provide information about acceptance rate of AI pair programmers.
+However, they do not provide information about the acceptance rate of AI pair programmers.
 Yetiştiren et al.
-(2023) perform a comparative study of GitHub Copilot, Amazon CodeWhisperer and ChatGPT (OpenAI, 2022) (sibling model of InstructGPT å(Ouyang et al., 2022)) in terms of code quality metrics like code correctness, code security, code reliability and code maintainability but do not consider acceptance of code by real users.
-A lot of studies on coding assistants do not measure their performance in terms of acceptance rate.
-Among those that do, we found that Ansible Lightspeed has the highest initial acceptance rate of 48.6%.
+(2023) perform a comparative study of GitHub Copilot, Amazon CodeWhisperer and ChatGPT (OpenAI, 2022) (sibling model of InstructGPT (Ouyang et al., 2022)) in terms of code quality metrics like code correctness, code security, code reliability and code maintainability but do not consider acceptance of code by real users.
 We did not find any user retention figures for any of the existing code completion systems or for the VS Code plugins in the VS Code marketplace.
-In the absence of more relevant baselines, we compare Lightspeed retention rate with that of Android and iOS apps.
 6. Limitations
@@
 We do not perform acceptance rate and edit analysis on all Ansible Lightspeed users.
-Our Telemetry data includes events from users that may not be Ansible programmers and simply want to try out the system; event data from these users would bias the acceptance rate and edit analysis.
-To mitigate this issue, we created a cohort of returning users, who used Ansible Lightspeed on two or more days.
+Our Telemetry data includes events from users who may not be Ansible programmers and simply want to try out the system with toy examples, which could add noise to our analysis.
+To mitigate this issue, we created a cohort of returning users who used Ansible Lightspeed on two or more days.
 Ansible Lightspeed’s acceptance rate is compared with models which are much more general, can be used for multiple programming languages and are likely bigger in terms of number of parameters and the amount of data used for training.
 These general models also tend to have a higher number of users overall and their performance on Ansible-YAML has not been studied or reported.
-One of the main purposes of our study was to explore the performance of a domain, language specific model vis-a-vis a more general model, with real users.
-The total number of users, 3,474, is sizable enough for us to be confident about our analysis.
+One of the main purposes of our study was to explore the performance of a domain, language-specific model vis-a-vis a more general model with real users.
+The total number of users, 3,910, is sizable enough for us to be confident about our analysis.
 Additionally, our evaluation of Ansible Lightspeed’s user retention lacked official data for comparisons with other code generation services.
-To address this gap, we compared its user retention with Android and iOS mobile apps, providing insights into its sustained user engagement.
-While the comparison methodology may not perfectly align with the domain, it highlights Ansible Lightspeed’s ability to retain users over time.
-Despite these considerations, the analysis and insights presented here offer a substantial understanding of Ansible Lightspeed’s effectiveness, user engagement, and potential areas for improvement.
+Lacking user retention baseline for code completion tools, we can’t say with confidence that users find Ansible Lightspeed useful enough to keep using it, especially for work.
+This is why we rely on temporal trends, acceptance rates, and explicit feedback to make such assessments.
+Despite these considerations, we hope the analysis and insights presented here offer a substantial understanding of Ansible Lightspeed’s effectiveness, user engagement, and potential areas for improvement.
 Based on this work, we hope to continue investigations into domain-specific code generation tools and their impact on IT automation.
@@
 (2023) show that a relatively small model fine-tuned on a specific Ansible task generation use case can outperform a much larger model on a benchmark dataset.
-In this paper we show that the suggestions made by such a specifically fine-tuned model can have relatively higher acceptance rate among a large set of users.
-We also show that the relatively higher acceptance rate results in better user retention compared to the average for Android and iOS apps and generally positive user feedback.
-Currently we support only Ansible task generation, however we are working to expand the capabilities of the model to be able to write entire Ansible playbook based on user prompt.
-We also plan to develop capabilities for Ansible code explanation, debugging and as well as customization for specific users.
+In this paper, by analyzing user interaction data of 3,910 users, we show that the suggestions made by such a specifically fine-tuned model can have a relatively higher acceptance rate among a large set of users.
+Analysis of feedback from 605 users and of 1,710 user edits to Lightspeed generated Ansible tasks led to interesting findings which we plan to use to improve Ansible Lightspeed.
+Currently, we use greedy decoding to generate one suggestion for every prompt.
+To account for user preferences and the fact that there can be multiple modules that can be used to implement similar functionalities, we hope to generate multiple suggestions.
+Also, currently we support only Ansible task generation, but we are working to expand the capabilities of the model to be able to generate longer Ansible sequences.
+With this, we hope to reduce the need for the user to reorganize the code and add missing details.
+We also plan to develop capabilities for Ansible code explanation, debugging, and as well as customization for specific users.
+We hope that our proposed user edit-based acceptance criteria are used by other code completion tools to get a better estimate of whether users truly found the model suggestion helpful.
+We also hope that by sharing Ansible Lightspeed’s user retention figures, others get encouraged to share similar figures for their code completion tools.
+And finally, by sharing the code for Ansible lightspeed service [^7] https://github.com/ansible/ansible-ai-connect-service and our analysis framework [^8] https://github.com/ansible-community/lightspeed-analysis-framework we hope to help others trying to develop similar tools.
 Acknowledgements.

### v4

@@
 Ansible is one such YAML-based IT automation-specific language.
-Ansible Lightspeed is an LLM-based service designed explicitly to generate Ansible YAML given natural language prompt.
-This paper first presents the design and implementation of the Ansible Lightspeed service.
-We then evaluate its utility to developers using diverse indicators, including extended utilization, analysis of user rejected suggestions, as well as analysis of user sentiments.
-The analysis is based on data collected for 10,696 real users including 3,910 returning users.
+Ansible Lightspeed is an LLM-based service designed explicitly to generate Ansible YAML, given natural language prompt.
+In this paper, we present the design and implementation of the Ansible Lightspeed service.
+We then evaluate its utility to developers using diverse indicators, including extended utilization, analysis of user edited suggestions, as well as user sentiments analysis.
+The evaluation is based on data collected for 10,696 real users including 3,910 returning users.
 The code for Ansible Lightspeed service and the analysis framework is made available for others to use.
-To our knowledge, our study is the first to involve thousands of users in evaluating code assistants for domain-specific languages.
-We propose an improved version of user acceptance rate and we are the first code completion tool to present N-Day user retention figures.
-With our findings we provide insights into the effectiveness of small, dedicated models in a domain-specific context.
-We hope this work serves as a reference for software engineering and machine learning researchers exploring code completion services for domain-specific languages in particular and programming languages in general.
-Large Language Models, Generative Models, Code Completion, IDE, User Study, Ansible
-$c$s: Computing methodologies Natural language processingccs: Software and its engineering Context specific languagesccs: Software and its engineering Application specific development environments
+[…]
+Large Language Models, Generative Models, Code Completion, IDE, User Study, Ansible, Acceptance Rate, IT Automation
 1. Introduction
@@
 Since it is open-source, we do not know the precise number of developers who use Ansible but we can get an indication of its popularity with the Ansible python community package (Team, 2024b) getting over 5 million downloads per month (Flynn, 2024).
-Ansible also has an active community on GitHub (Team, 2024a) with about 5,000 contributors and about 54,000 commits on a yearly basis, which is likely a small fraction of the actual user base.
+Ansible also has an active community on GitHub (Team, 2024a) with about 5,000 contributors and about 54,000 commits on an yearly basis, which is likely a small fraction of the actual user base.
 [^1] Another indication of Ansible’s popularity is from Ansible’s social media accounts of Reddit (Team, 2024c) and X (Team, 2024d), counting 63,000 readers and 72,000 followers, respectively.
@@
 (2023) showed that a relatively small model fine-tuned on high quality Ansible data can outperform a much larger and more general model on benchmark Natural Language to Ansible-task generation test data.
-But this work does not involve real users in their development environment interacting with a model.
+But this work does not involve real users in their development environment.
 Red Hat Ansible Lightspeed (RedHat, 2023), further referred to as Ansible Lightspeed or simply Lightspeed, is a generative AI service that utilizes watsonx Code Assistant for Ansible (WCA), an extension of Ansible Wisdom (Pujar et al., 2023), to produce code recommendations based on best practices.
@@
   • We are the first code completion tool to share user retention statistics. This can be used as a baseline for similar tools in the future.
-  • We propose an improved acceptance criteria that checks if the accepted suggestion was actually used by the user.
+  • We propose an improved acceptance criteria, called Strong Acceptance rate, that checks if the accepted suggestion was actually used by the user.
   • We show that Ansible Lightspeed’s acceptance rate is higher than that of other, more general, code completion tools.
@@
 The model is pre-trained with about 143 billion tokens and then fine-tuned with about 500k natural language to Ansible task samples.
-The fine-tuning data is created using a technique similar to the one of Ansible Wisdom (Pujar et al., 2023).
+The training environment, hyper-paramters and fine-tuning data creation techniques are similar to Ansible Wisdom (Pujar et al., 2023).
 The pre-training and fine-tuning data is license filtered to remove all data samples coming from files or repositories with restrictive licenses.
@@
 The pre-training data is also deduplicated at a file level.
-The model server runs on machines containing A100100100100 GPUs equipped with sufficient memory to return an inference response within $22$ seconds on average.
+The model server runs on machines containing A100100100100 GPUs equipped with sufficient memory to restrict the end-to-end latency within $22$ seconds on average.
+Every model suggestion is a multi-line Ansible task with 6.13 lines and 20.1 tokens on average.
 Post-processing
@@
 It also redacts any credentials and removes and/or replaces any deprecated Ansible syntax from the suggestions.
-After this post-processing stage is complete, the response is sent back to the client.
+After post-processing is complete, the response is sent back to the client.
 2.3. Content Matching Pipeline
@@
 4.
-An event tracking platform monitors and records a wide range of events about user interactions, and system activities.
+The event tracking platform monitors and records a wide range of events about user interactions, and system activities.
 This platform is a critical part of our data collection pipeline, ensuring that we systematically capture important events within our system.
@@
 3.1. Temporal Analysis
-Lightspeed’s goal is to improve the productivity of developers who use Ansible for work.
-Understanding when requests were made could give some hint that Lightspeed is being used for work related queries.
+Lightspeed’s goal is to improve the productivity of developers who use Ansible for work. Understanding when requests were made could give hint that Lightspeed is being used for work.
 Figure 6 shows all Lightspeed completion requests [^4] The figures here represent telemetry events, which may contain duplicates.
@@
 However, we need more analysis to be sure.
-In Figure 6 we can also see a significant decline in the completion requests between December 23, 2023, and January 12, 2024.
-This could be due to the Holiday season when many offices are usually closed.
+In Figure 6 we can also see a significant decline in the completion requests between December 23, 2023, and January 12, 2024. This could be due to the Holiday season when many offices are closed.
 Q: When do users use Ansible Lightspeed?
@@
 Total suggestions 62,099
+Average number of lines per suggestion 6.13
+Average number of tokens per suggestion 20.1
 Initially accepted suggestions 40,938
@@
 In 24,811 instances (60.60% of the accepted suggestions), Ansible Lightspeed’s suggestions were used without edits so we can assume that they aligned with user requirements.
-In 5,672 scenarios (13.85% of the accepted suggestions), users made small adjustments, less than 50% edits [^6] Suggestion analyzer part of Sec.
-2.4, which explains how edits are calculated., to the accepted suggestions.
+In 5,672 scenarios (13.85% of the accepted suggestions), users made small adjustments, less than 50% edits, [^6] Suggestion analyzer part of Sec.
+2.4 explains how edits are calculated.
+to the accepted suggestions.
 An example of such a light or minor edit would be, when Ansible Lightspeed accurately predicts the correct module, users modified one or more module keys (options) and/or values to better align with their specific needs.
@@
 One explanation is that users found it challenging to review the suggestion in the greyed text (inline suggestion text) within the VS Code editor, leading them to initially accept and then delete after closer review.
-Another possible reason could be that they were experimenting with different prompts trying one after another and discarding the answers.
+Another reason could be that they were experimenting with different prompts trying one after another and discarding the answers.
 We further analyzed 5,672 instances which were lightly edited, to make sure that they were indeed minor edits. Figure 8.b gives a break down of different kinds of minor edits.
@@
 The edit analysis described above shows that the acceptance rate as currently defined can be misleading.
-Which is why we define a new acceptance rate metric, called Edit¡50% acceptance rate.
-Edit¡50% Acceptance rate is the percentage of accepted suggestions, in which we consider a suggestion accepted only if the following conditions are fulfilled:
+Which is why we define a new acceptance rate metric, called Strong Acceptance rate.
+Strong Acceptance rate is the percentage of accepted suggestions, where a suggestion is considered accepted only if:
   1. (1)
@@
     Critical components, like the module name in case of Ansible, should not be edited.
-As per this definition, the Edit¡50% acceptance rate for Ansible Lightspeed comes out to 49.08%, compared to the initial acceptance rate of 65.92%.
-We feel Edit¡50% acceptance rate more closely aligns with users perception of what they consider accepted, compared to the initial acceptance rate.
+As per this definition, the Strong Acceptance rate for Ansible Lightspeed comes out to 49.08%, compared to the initial acceptance rate of 65.92%.
+We feel Strong Acceptance rate more closely aligns with users perception of what they consider accepted, compared to the initial acceptance rate.
 Table.
@@
 It does not give any insight into whether the user accepted the suggestion as is, or edited the suggestion or altogether deleted it.
-Ansible Lightspeed generates multi-line Ansible code suggestions. Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and Edit¡50% acceptance rate.
+Ansible Lightspeed always generates multi-line Ansible code suggestions, with an average of 6.13 lines and 20.1 tokens per suggestion over all the suggestions shown in Table.
+1.
+Lightspeed has a higher acceptance rate compared to the other tools for both initial accept and Strong Acceptance rate.
 Table 2.
@@
 Ansible Lightspeed Initial accept are percent suggestions which were accepted by users.
-Edit¡50% are percent suggestions which were accepted and had less than 50% user edits, based on difflib SequenceMatcher(Library, 2001).
+Strong Acceptance rate is percent suggestions which were accepted and had less than 50% user edits, based on difflib SequenceMatcher(Library, 2001).
 Service Acceptance Rates User Count
@@
 Initial accept (Ansible) 65.9% 3,910
-Edit¡50% (Ansible) 49.1% 3,910
+Strong accept (Ansible) 49.1% 3,910
 Q: At what rate do users accept suggestions?
@@
 65.9% of suggestions are initially accepted.
-We define new acceptance rate, Edit¡50% Acceptance rate.
-49.1% of model suggestions were accepted as per Edit¡50% Acceptance rate.
+We define new acceptance rate, Strong Acceptance rate.
+49.1% of model suggestions were accepted as per Strong Acceptance rate.
 Ansible Lightspeed has relatively higher acceptance for Ansible, compared to more general tools for various programming languages.
-3.5. Rejected Suggestion Analysis
+3.5. Edited Module Analysis
 There are a total of 1,710 initially accepted suggestions which were explicitly edited by the users at the module level.
-We consider such suggestions to be rejected for Edit¡50% Acceptance rate.
+We consider such suggestions to be rejected for Strong Acceptance rate.
 However, the users must have found something useful in the suggestion for them to accept it and then spend sometime editing it.
 We analyse these edits in order to better understand how the model suggestions differ from user expectations.
-3.5.1. User preference
+The edit categories were identified based on manual analysis.
+The number of samples in each category was obtained based on rules to find similar samples.
+A suggestion could have been edited for multiple reasons, so an edited sample can belong to multiple categories.
+User preference
 In order to remove ambiguity in module names, Ansible Lint encourages users to use Fully Qualified Collection Name, or FQCN (Sbarnea, 2023), when declaring modules.
@@
 The body of the generated Ansible task often had little to no edits, indicating that the FQCN module name, which is correct, is not what they prefer.
-3.5.2. YAML reorganization
+YAML reorganization
 As per our observation, in the process of writing an Ansible playbook users often reorganize the YAML structure.
@@
 There are at least 211 (12.3%) such instances of reorganization.
-3.5.3. Functionally similar modules
+Functionally similar modules
 Module confusion usually occurs when there are multiple functionally similar modules which can be used to implement user’s intent and the user is at least somewhat familiar with one of them.
@@
 Apart from command or shell module related edits, we found 336 (19.6%) instances of edits related to modules which have similar functionalities and can be used interchangeably.
-There are another 586 (28%) instances where the user edited the generated module. Understanding these edits requires further analysis.
+There are another 586 (34.2%) cases where the user edited the module. Understanding these edits requires more analysis.
 Q: Why do users edit module names in accepted suggestions?
@@
 4. User Feedback
-A total of 605 users have provided rating and comments using the the sidebar for Ansible Lightspeed in the VS Code Editor, as shown in Figure 4.
+A total of 605 users have provided rating and comments using the the sidebar for Ansible Lightspeed in the VS Code Editor (Figure 4).
 4.1. Star rating
@@
 Apart from star-rating, users also submit feedback in the form of textual comments.
-By analyzing these comments we hope to find areas for improvement and also know what users like about the service.
-We manually review and classify negative feedback (1 or 2 stars) and positive feedback (4 or 5 stars) into various reasons.
+We manually review and classify negative feedback (1 or 2 stars) and positive feedback (4 or 5 stars).
+The categories are discovered in the data using open coding.
+Each comment is assigned to only one category, the main point of the feedback.
 We also share some examples of real user comments, both positive and negative.
@@
 For instance, they may ask for multi-task or playbook generation, but the model is currently only trained for generating single tasks:
-“Has really bad recommendations. Tried creating a playbook with name ”Get …”” and it get me a module recommendation”
-Finally, 5.79% of users are dissatisfied with their overall experience, specifically citing the need to re-authenticate periodically as an issue. The remaining 12.04% of comments are not informative.
+“Has really bad recommendations. Tried creating a playbook with name ”Get …” and it get me a module recommendation”
+Finally, 5.79% of users are dissatisfied with their overall experience, specifically citing the need to re-authenticate periodically as an issue.
+The remaining 12.04% of comments are not informative.
 Positive feedback
@@
 Breakdown of reasons for positive experience is presented in Figure 11.
-Most users, about 42.7%, consider improved productivity as Ansible Lightspeed’s key benefit.
+About 42.7% of users consider improved productivity as Ansible Lightspeed’s key benefit.
 Multiple users mentioned that they noticeably saved time by using Lightspeed.
@@
 We also plan to develop capabilities for Ansible code explanation, debugging, and as well as customization for specific users.
-We hope that our proposed user edit-based acceptance criteria are used by other code completion tools to get a better estimate of whether users truly found the model suggestion helpful.
+We hope that our proposed user edit-based acceptance criteria, Strong Acceptance rate, is used by other code completion tools to get a better estimate of whether users truly found the model suggestion helpful.
 We also hope that by sharing Ansible Lightspeed’s user retention figures, others get encouraged to share similar figures for their code completion tools.